# SpectrumWorld: Artificial Intelligence Foundation for Spectroscopy

### v3

@@ Appendix A SpectrumBench Detailed Information
Appendix B SpectrumAnnotator Technical Details
@@
MolPuzzle (2024b) represents the first benchmark specifically designed for LLMs in spectroscopic analysis, employing a three-stage approach to generate question-answer pairs. While this template-based generation method offers efficiency, it suffers from limited coverage of spectroscopic domains and overly simplistic question formats. In the field of spectroscopy, high-quality data and benchmarks are crucial to advance AI research. The design of SpectrumAnnotator originates from two key insights: First, the process of creating benchmarks shares similarities with the supervised data generation methods used in LLM pre-training and post-processing. Just as high-quality training data is essential for model performance, well-designed benchmarks are equally critical for evaluating and advancing the field. Second, we aim to utilize LLMs’ few-shot and zero-shot capabilities to generate diverse benchmarks, enabling batch processing of seed datasets to construct large-scale pre-training and post-processing data. Additionally, we leverage LLMs’ discriminative abilities for preliminary data screening and establish closed-loop mechanisms for continuous improvement.
@@
Appendix C Benchmarking Candidates
Open-source Models
@@ Appendix C Benchmarking Candidates
Qwen2.5-VL-32B-Instruct(2025). Alibaba’s open-source Vision-Language multimodal large model that handles reasoning and generation for images, text and video. It employs a hierarchical tagging architecture, supports multi-turn conversations and complex reasoning, and both the model weights and code are publicly available.
@@
Qwen2.5-VL-72B-Instruct(2025). Qwen2.5’s larger-scale model enhances cross-modal reasoning and instruction-following capabilities, delivering superior performance on benchmarks such as MMMU and M3Exam while supporting multitasking and multilingual inputs - and is completely open-source.
@@
InternVL3-78B (2024b). Shanghai AI Lab releases the multimodal model, combining native multimodal pre-training, variable visual position encoding (V2PE), MPO, and test-time scaling to approach GPT-4o performance.
@@
Llama-3.2-11b-Vision-Instruct(2024). Meta’s 11 B lightweight multimodal model locks Llama-3.1 8 B text and pairs it with a ViT encoder. Two-stage training: image-text alignment then SFT+DPO, using RoPE-2D. Open-source.
@@
Lllama-3.2-90b-Vision-Instruct(2024). The 90B features a more advanced vision adapter with cross-attention layers to inject image features into the LLM core. It is tuned with SFT and RLHF for enhanced performance on complex visual reasoning tasks.
@@
DeepSeek-VL-2(2024). An open-source model from DeepSeek-AI featuring a Mixture-of-Experts (MoE) backbone and a dynamic tiling vision encoder for high-resolution images. It achieves or exceeds the state-of-the-art performance at the time on benchmarks like MMMU and DocVQA, with its code and weights fully available on GitHub.
@@
Doubao-1.5-Vision-Pro (2025). It features a dynamic resolution visual encoder and MoE architecture, supporting visual QA, text-image matching, and image description. With billions of parameters, it shows strong generalization across scenarios and is available for self-hosting and fine-tuning.
@@
Doubao-1.5-Vision-Pro-Thinking (2025). It integrates a “Deep Thinking Mode” and is trained with multi-round Reward Learning and reasoning style training. It excels in scientific, mathematical, and chain-of-thought reasoning. Supports open-source calling and API integration.
Closed-source Models
@@
GPT-4o (2024). OpenAI’s flagship “omni” model natively supports text, audio, and image modalities. Delivers GPT-4-level intelligence with significantly faster response times and enhanced multimodal capabilities.
@@
GPT-4.1(2025). A reinforced version of GPT-4 deployed through the OpenAI API, offering improved handling of complex instructions and logical reasoning; accepts multimodal inputs but is primarily geared toward text-centric tasks.
@@
GPT-4-Vision(2023). A version of GPT-4 equipped with image input capabilities, optimized for understanding images and text and for the generation of conversational content, widely used for image-based Q&A.
@@
Claude-3.5-Sonnet (2024). Anthropic’s multimodal large language model has mixed inference capabilities and powerful visual understanding functions. It supports a context of 200K tokens and is skilled in natural writing and code generation.
Claude-3.7-Sonnet (2025a). An evolution of Claude3.5 Sonnet that introduces hybrid reasoning—users can choose between fast modes or step‑by‑step logical chains; offers strong task flexibility, extended context windows, and deep instruction‑following in multimodal settings.
Claude-4-Opus (2025b). Anthropic’s flagship model, designed for complex tasks. It boasts a powerful memory architecture and parallel tool invocation capabilities, and integrates with Claude Code, performing exceptionally in coding and reasoning benchmark tests.
Claude-4-Sonnet (2025b). Claude-3.7-Sonnet’s successor, balancing performance and speed, with low latency and high resource efficiency, excels in code generation.
Grok-2-Vision(2024). The multi-modal model of xAI combines language and visual processing capabilities to handle various images and documents, and supports multilingual recognition and style analysis.
Qwen-VL-Max(2024). The closed-source flagship model of Alibaba’s Qwen series has been optimized for deployment in enterprise-level multimodal tasks, supporting joint input of images, text, videos, and others, with ultra-large parameter volume and high inference capability.
Appendix D Error Analysis
Signal Level
@@ Appendix D Error Analysis
Appendix E Model Accuracy vs. Token Assumptions
@@
Appendix F Detailed Data Structure
This section provides a rigorous overview of the three principal data structures that underpin our work: seed datasets, benchmark data, and evaluation results.
Seed Datasets Structure
@@ Appendix F Detailed Data Structure
1 {
2 "molecule_index": "MOL_0001",
3 "smiles": "CCCCC1=CC=CC=C1",
4 "formula": "C10H14",
5 "spectra": [
6 {"spectrum_type": "IR", "path": ["IR/MOL_0001.png"]},
7 {"spectrum_type": "MASS", "path": ["MASS/MOL_0001.jpg", "MASS/MOL_0001_2.jpg"]},
8 {"spectrum_type": "C-NMR", "path": ["C-NMR/MOL_0001.png"]},
9 {"spectrum_type": "H-NMR", "path": ["H-NMR/MOL_0001.png"]}
10 ]
11 }
SpectrumBench Data Structure
@@
1 {
2 "id": "Perception_a9cf_250723_235951_318294",
3 "image_path": [
4 "data/Perception/Basic Property Prediction/Perception_a9cf_q.png"
5 ],
6 "question": "Given the mass spectrum image, what is the most likely molecular ion peak (m/z) observed for this compound?",
7 "choices": ["85", "90", "120", "133"],
8 "answer": "133",
9 "category": "Perception",
10 "sub_category": "Basic Property Prediction",
11 "source": "",
12 "timestamp": "2025-07-23 23:59:51"
13 }
Evaluation Results Structure
@@
1 {
2 "id": "Signal_9131_250723_110552_245529_2",
3 "image_path": [
4 "data/Signal/Spectrum Type Classification/Signal_9131_2_q.png"
5 ],
6 "question": "What type of spectrum is shown in the image?",
7 "choices": [
8 "Infrared Spectrum (IR)",
9 "Proton Nuclear Magnetic Resonance (H-NMR)",
10 "Mass Spectrometry (MS)",
11 "Carbon Nuclear Magnetic Resonance (C-NMR)"
12 ],
13 "answer": "Mass Spectrometry (MS)",
14 "category": "Signal",
15 "sub_category": "Spectrum Type Classification",
16 "source": "",
17 "timestamp": "2025-07-23 11:05:52",
18 "model_prediction": "Mass Spectrometry (MS)",
@@
20 "pass": true
21 }
Appendix G Cost Analysis
@@
Table 4 summarizes the key statistics for representative models evaluated in this study. Detailed experimental tracking information can be found in our anonymous link at https://ai4s-chem.github.io/SpectrumWorld/.
@@
Appendix H Limitations
@@
Limitations of SpectrumBench First, regarding Task Format, SpectrumBench currently supports only multiple-choice and a limited number of open-ended questions. While this design is suitable for Large Language Models (LLMs), it is insufficient for evaluating a broader range of machine learning models, such as Convolutional Neural Networks (CNNs) and Graph Neural Networks (GNNs), as discussed in our introduction. Second, concerning Spectrum Type, although we have incorporated a wide array of spectrum types compared to previous works(2025; 2025; 2024a; 2024a), several crucial spectroscopic modalities remain uncovered. Notable examples include X-ray Diffraction (XRD) (2024a; 2023) and fluorescence spectra (1962), which are vital for comprehensive material characterization. Finally, addressing Spectroscopic Task Type, spectroscopy techniques are fundamental across diverse scientific disciplines, including physics, astronomy, chemistry, and biology, primarily for characterizing substances like molecules, proteins, peptides, and SMILES sequences. From the perspective of LLMs, a generic categorization of modalities into “text” and “images” is inadequate for representing the complexity of data. The inherent diversity of spectroscopic modalities complicates the immediate definition of all possible tasks. Consequently, SpectrumBench presently lacks important benchmarks in several areas, such as spectrum-spectrum retrieval (1969; 2022; 2025) and peptide sequence analysis (2024a). We acknowledge that it will be challenging for SpectrumBench to encompass all relevant tasks in the near future, and we aim to foster collaborative efforts with the community and various laboratories to collectively advance the development of AI in spectroscopy.

### v4

@@ B.5 Data Distribution
Appendix C SpectrumAnnotator Technical Details
@@
MolPuzzle (Guo et al., 2024b) represents the first benchmark specifically designed for LLMs in spectroscopic analysis, employing a three-stage approach to generate question-answer pairs. While this template-based generation method offers efficiency, it suffers from limited coverage of spectroscopic domains and overly simplistic question formats. In the field of spectroscopy, high-quality data and benchmarks are crucial to advance AI research. The design of SpectrumAnnotator originates from two key insights: First, the process of creating benchmarks shares similarities with the supervised data generation methods used in LLM pre-training and post-processing. Just as high-quality training data is essential for model performance, well-designed benchmarks are equally critical for evaluating and advancing the field. Second, we aim to utilize LLMs’ few-shot and zero-shot capabilities to generate diverse benchmarks, enabling batch processing of seed datasets to construct large-scale pre-training and post-processing data. Additionally, we leverage LLMs’ discriminative abilities for preliminary data screening and establish closed-loop mechanisms for continuous improvement.
@@
Appendix D Benchmarking Candidates
Figure 14: Performance ranking of various LLMs.
D.1 Open-source Models
@@
Qwen2.5-VL-32B-Instruct(Bai et al., 2025). Alibaba’s open-source Vision-Language multimodal large model that handles reasoning and generation for images, text and video. It employs a hierarchical tagging architecture, supports multi-turn conversations and complex reasoning, and both the model weights and code are publicly available.
@@
Qwen2.5-VL-72B-Instruct(Bai et al., 2025). Qwen2.5’s larger-scale model enhances cross-modal reasoning and instruction-following capabilities, delivering superior performance on benchmarks such as MMMU and M3Exam while supporting multitasking and multilingual inputs - and is completely open-source.
@@
InternVL3-78B (Chen et al., 2024b). Shanghai AI Lab releases the multimodal model, combining native multimodal pre-training, variable visual position encoding (V2PE), MPO, and test-time scaling to approach GPT-4o performance.
@@
Llama-3.2-11b-Vision-Instruct(Meta AI, 2024). Meta’s 11 B lightweight multimodal model locks Llama-3.1 8 B text and pairs it with a ViT encoder. Two-stage training: image-text alignment then SFT+DPO, using RoPE-2D. Open-source.
@@
Lllama-3.2-90b-Vision-Instruct(Meta AI, 2024). The 90B features a more advanced vision adapter with cross-attention layers to inject image features into the LLM core. It is tuned with SFT and RLHF for enhanced performance on complex visual reasoning tasks.
@@
DeepSeek-VL-2(Wu et al., 2024). An open-source model from DeepSeek-AI featuring a Mixture-of-Experts (MoE) backbone and a dynamic tiling vision encoder for high-resolution images. It achieves or exceeds the state-of-the-art performance at the time on benchmarks like MMMU and DocVQA, with its code and weights fully available on GitHub.
@@
Doubao-1.5-Vision-Pro (Doubao Team, 2025). It features a dynamic resolution visual encoder and MoE architecture, supporting visual QA, text-image matching, and image description. With billions of parameters, it shows strong generalization across scenarios and is available for self-hosting and fine-tuning.
@@
Doubao-1.5-Vision-Pro-Thinking (Doubao Team, 2025). It integrates a “Deep Thinking Mode” and is trained with multi-round Reward Learning and reasoning style training. It excels in scientific, mathematical, and chain-of-thought reasoning. Supports open-source calling and API integration.
GLM-4.5V(VTeam, 2025). An open-source vision-language model from Zhipu AI and Tsinghua University that introduces a versatile ”thinking paradigm” for enhanced reasoning. It leverages scalable reinforcement learning and supports full-spectrum vision reasoning, including GUI agent operations and code generation from screenshots.
InternS1(Intern-S1 Team, 2025).A vision-language model developed by Shanghai AI Laboratory that features a specialized ”Thinking” mode for enhanced multi-step reasoning. This mode allows the model to perform a series of self-guided logical steps to solve complex problems, particularly in scientific, mathematical, and logical domains.
D.2 Closed-source Models
@@
GPT-4o (OpenAI, 2024). OpenAI’s flagship “omni” model natively supports text, audio, and image modalities. Delivers GPT-4-level intelligence with significantly faster response times and enhanced multimodal capabilities.
@@
GPT-4.1(OpenAI, 2025). A reinforced version of GPT-4 deployed through the OpenAI API, offering improved handling of complex instructions and logical reasoning; accepts multimodal inputs but is primarily geared toward text-centric tasks.
@@
GPT-4-Vision(OpenAI, 2023). A version of GPT-4 equipped with image input capabilities, optimized for understanding images and text and for the generation of conversational content, widely used for image-based Q&A.
@@
Claude-3.5-Sonnet (Anthropic, 2024). Anthropic’s multimodal large language model has mixed inference capabilities and powerful visual understanding functions. It supports a context of 200K tokens and is skilled in natural writing and code generation.
Claude-3.7-Sonnet (Anthropic, 2025a). An evolution of Claude3.5 Sonnet that introduces hybrid reasoning—users can choose between fast modes or step‑by‑step logical chains; offers strong task flexibility, extended context windows, and deep instruction‑following in multimodal settings.
Claude-4-Opus (Anthropic, 2025b). Anthropic’s flagship model, designed for complex tasks. It boasts a powerful memory architecture and parallel tool invocation capabilities, and integrates with Claude Code, performing exceptionally in coding and reasoning benchmark tests.
Claude-4-Sonnet (Anthropic, 2025b). Claude-3.7-Sonnet’s successor, balancing performance and speed, with low latency and high resource efficiency, excels in code generation.
Grok-2-Vision(xAI, 2024). The multi-modal model of xAI combines language and visual processing capabilities to handle various images and documents, and supports multilingual recognition and style analysis.
Qwen-VL-Max. The closed-source flagship model of Alibaba’s Qwen series has been optimized for deployment in enterprise-level multimodal tasks, supporting joint input of images, text, videos, and others, with ultra-large parameter volume and high inference capability.
Gemini-2.5-Pro(Gemini Team, ).A multimodal model from Google DeepMind that achieves state-of-the-art performance on frontier reasoning and coding benchmarks. It excels at multimodal understanding, including the ability to process up to 3 hours of video content and convert it into interactive code. Its combination of long context, multimodality, and enhanced reasoning capabilities unlocks new agentic workflows and complex problem-solving.
Appendix E Error Cases Study
E.1 Signal Level
@@ E.4 Generation Level
Appendix F Model Accuracy vs. Token Assumptions
@@
Appendix G Detailed Data Structure
This section details the comprehensive seed datasets curation pipeline and the three primary data structures that underpin our framework: the foundational seed datasets, the structured benchmark data, and the standardized evaluation results.
G.1 Seed Data Curation Details
The seed datasets are curated from three primary sources to ensure both diversity and scientific rigor:
  1. 1.
    Proprietary collections and in-house experimental data: These include unpublished spectroscopic measurements and curated datasets from our collaborating laboratories. This source comprises approximately 238,869 molecular data points covering 8 types of spectra, offering higher authenticity and usability compared to most computationally generated spectra.
  2. 2.
    Public repositories and benchmark datasets: We integrate data from a range of widely recognized and authoritative sources, including SDBS (of Advanced Industrial Science &, AIST), QM9S (Zou et al., 2023), NovoBench (Zhou et al., 2024a), and MolPuzzle (Guo et al., 2024b), among others. In total, seven distinct repositories and public datasets are used, collectively encompassing over 1.01 million unique chemical compounds.
  3. 3.
    Literature mining: Spectral data are systematically extracted from the Supporting Information sections of peer-reviewed publications, with a focus on articles from leading journals such as the Journal of the American Chemical Society (JACS) and ACS Catalysis.
All collected datasets undergo a unified processing pipeline that systematically maps each entry into three core chemical spaces: SMILES, molecular formula, and spectra. The resulting seed datasets are organized at the level of individual chemical substances, with each record containing the compound’s SMILES, molecular formula, and a structured set of associated spectra, all stored in a standardized JSON format. This robust foundation facilitates downstream annotation and interoperability.
G.2 Seed Datasets Structure
@@
"molecule_index": "MOL_0001",
"smiles": "CCCCC1=CC=CC=C1",
"formula": "C10H14",
"spectra": [
{"spectrum_type": "IR", "path": ["IR/MOL_0001.png"]},
{"spectrum_type": "MASS", "path": ["MASS/MOL_0001.jpg",
"MASS/MOL_0001_2.jpg"]},
{"spectrum_type": "C-NMR", "path": ["C-NMR/MOL_0001.png"]},
{"spectrum_type": "H-NMR", "path": ["H-NMR/MOL_0001.png"]}
G.3 SpectrumBench Data Structure
@@
"id": "Perception_a9cf_250723_235951_318294",
"image_path": [
"data/Perception/Basic Property Prediction/Perception_a9cf_q.png"
"question": "Given the mass spectrum image, what is the most likely
molecular ion peak (m/z) observed for this compound?",
"choices": ["85", "90", "120", "133"],
"answer": "133",
"category": "Perception",
"sub_category": "Basic Property Prediction",
"source": "",
"timestamp": "2025-07-23 23:59:51"
G.4 Evaluation Results Structure
@@
"id": "Signal_9131_250723_110552_245529_2",
"image_path": [
"data/Signal/Spectrum Type Classification/Signal_9131_2_q.png"
"question": "What type of spectrum is shown in the image?",
"choices": [
"Infrared Spectrum (IR)",
"Proton Nuclear Magnetic Resonance (H-NMR)",
"Mass Spectrometry (MS)",
"Carbon Nuclear Magnetic Resonance (C-NMR)"
"answer": "Mass Spectrometry (MS)",
"category": "Signal",
"sub_category": "Spectrum Type Classification",
"source": "",
"timestamp": "2025-07-23 11:05:52",
"model_prediction": "Mass Spectrometry (MS)",
@@
"pass": true
Appendix H Cost Analysis
@@
Table 7 summarizes the key statistics for representative models evaluated in this study.
@@
Appendix I Usage of Large Language Models in This Manuscript
In preparing this manuscript, we used a large language model (LLM) solely for editorial purposes. Its functions were limited to proofreading for typographical errors, correcting grammatical mistakes, and enhancing the clarity and readability of the text.
Appendix J Limitations
@@ Appendix I Usage of Large Language Models in This Manuscript
Limitations of SpectrumBench First, regarding Task Format, SpectrumBench currently supports only multiple-choice and a limited number of open-ended questions. While this design is suitable for Large Language Models (LLMs), it is insufficient for evaluating a broader range of machine learning models, such as Convolutional Neural Networks (CNNs) and Graph Neural Networks (GNNs), as discussed in our introduction. Second, concerning Spectrum Type, although we have incorporated a wide array of spectrum types compared to previous works (Lu et al., 2025; Xu et al., 2025; Bushuiev et al., 2024b; Zhou et al., 2024a), several crucial spectroscopic modalities remain uncovered. Notable examples include X-ray Diffraction (XRD) (Guo et al., 2024a; Salgado et al., 2023) and fluorescence spectra (Parker & Rees, 1962), which are vital for comprehensive material characterization. Finally, addressing Spectroscopic Task Type, spectroscopy techniques are fundamental across diverse scientific disciplines, including physics, astronomy, chemistry, and biology, primarily for characterizing substances like molecules, proteins, peptides, and SMILES sequences. From the perspective of LLMs, a generic categorization of modalities into “text” and “images” is inadequate for representing the complexity of data. The inherent diversity of spectroscopic modalities complicates the immediate definition of all possible tasks. Consequently, SpectrumBench presently lacks important benchmarks in several areas, such as spectrum-spectrum retrieval (Curry et al., 1969; Wang et al., 2022; Lu et al., 2025) and peptide sequence analysis (Zhou et al., 2024a). We acknowledge that it will be challenging for SpectrumBench to encompass all relevant tasks in the near future, and we aim to foster collaborative efforts with the community and various laboratories to collectively advance the development of AI in spectroscopy.